OpenStudy (anonymous):

kelli lights a candle after 4 minutes it's 8 inches tall after 16 minutes it's 5 inches tall which equation in slope intercept form represents the height of the candle after x minutes y=1/4x+9, y=1/4x+4, y=4x+24, or y=4x+4

OpenStudy (chris):

ok, so slope intercept form is always y = mx +b - where m is the rate of change, and b is the constant

OpenStudy (chris):

in this problem, they're giving you the "y" value - the height of the candle

OpenStudy (chris):

so I"m going to make you work through this one, but it will be worth it

OpenStudy (chris):

so they give you two equations with the two pieces of information in the word problem

OpenStudy (chris):

in both cases, they give you y - the height of the candle, and x - the time since it was lit

OpenStudy (chris):

1. 8 = m*4 +b - where m is the "slope" - or rate of change of the height of the candle 2. 5 = m*16 + b

OpenStudy (chris):

can you solve for the two variables given those two equations?

OpenStudy (chris):

let me know, brb

OpenStudy (anonymous):

well srry i will try to do it i have to get ready for bed now